# Managing Classical Processing Requirements for Quantum Error Correction

### v1

@@ 3.3. Concurrency and Delayed Decoding
Critical decodes must be serviced during the execution of a program to avoid the increase in memory requirements and processing times discussed above. If there are many concurrent critical decodes that occur frequently during the execution of a program, an appropriate number of decoders will be needed to process these critical decodes. The level of concurrency depends entirely on the layout used to build the quantum computer – layouts such as the Compact (Litinski2019, , p. 7), Fast (Litinski2019, , p. 9), and the Edge-Disjoint Path (EDPC) (Beverland2022edpc, ) are some proposed layouts that can be used to build fault-tolerant quantum computers using the Surface Code. Layouts determine the number of physical qubits required – for example, the Compact layout requires the fewest physical qubits since there is a single routing lane at the cost of completely serializing operations. The Fast and EDPC layouts allow more concurrency at the cost of more physical qubits.
What is the average level of concurrency when executing a quantum program on a fault-tolerant quantum computer? Fig. 4(e) shows a histogram of the number of critical decodes [^3] For the Surface Code, we assume every logical qubit requires two decoders – one each for the $X$ and $Z$ observables. for select workloads generated by the Lattice Surgery Compiler (watkins2023high, ). This histogram shows that the peak concurrency is attained very infrequently. This implies that most logical qubits function as memory qubits or execute Clifford gates more often than $T$-gates, and this can allow some qubits to not be decoded in real-time.
Quantum programs are serial in terms of $T$-gates applied – not every qubit always requires access to a fast decoder.
@@ 3.4. Goal: Make Classical Processing Efficient
Provisioning a hardware decoder for every logical qubit in the system can be resource intensive – Fig. 4(f) shows an estimate of the number of FPGAs required just for decoding when (i) a single distillation factory is used, and (ii) an optimal number of distillation factories are used for different workloads ($\sim$10FPGA LUTs/decoder (Vittal2023, ; Riverlane2023, )). This estimation was done using the Azure QRE (Beverland2022, ) that uses the Fast layout. Note that the total hardware requirement will be significantly higher because of control and readout components. Having shown how the syndrome processing rate is crucial in ensuring that computation does not require excessive memory and time and how quantum programs are inherently serial in terms of critical decodes, the question we seek to answer in this work is–
How can we minimize the use of hardware decoders and lower classical processing costs without sacrificing the performance and reliability?
4. Virtual Quantum Decoders
We now show how the number of hardware decoders can be reduced to be less than the number of logical qubits in the system, and how decoding can be scheduled to prevent excessive accumulation of undecoded syndromes.
@@ 4. Virtual Quantum Decoders
4.1. Working with Fewer Hardware Decoders
Reducing the number of available decoders implies that qubits will share hardware resources, resulting in time-division multiplexing of decoder instances among logical qubits. Fig. 5 shows how compared to a system with decoders for every logical qubit in the system ($N$ qubits, $N$ decoders), a system with fewer ($M$, $N>M$) decoders will require resources to be shared with time.
Before we discuss different decoder scheduling policies, it is important to understand the time granularity at which any scheduling policy will operate on. Since logical operations (Clifford or non-Clifford) in a Surface Code error-corrected quantum computer will require at least $d$ rounds before the next operations, we define a slice (watkins2023high, ) as the smallest time step between logical operations that a decoder scheduling policy can work on. Every slice consists of $d$ rounds of syndrome measurements, thus making the scheduling policy agnostic of the actual code distance used.
@@ 4.1. Working with Fewer Hardware Decoders
4.2. Decoder Scheduling Policies
Static decoder scheduling refers to decoder scheduling that can be performed at compile time. Since most quantum programs do not have any control-flow instructions, scheduling can be performed statically. Scheduling decoders is similar to CPU scheduling performed by all operating systems today, where the number of processes is more than the number of available processor cores (Liu1973, ).
Longest Undecoded Sequence: To quantify the fairness of a decoder scheduling policy, we use ‘Longest Undecoded Sequence’, which measures how well the decoders are servicing all logical qubits. A large undecoded sequence length implies that a qubit has been left undecoded for a long time – increasing the memory consumed to store undecoded syndromes. Fig. 6(a) shows an example of determining the longest undecoded sequence length.
Consider an arbitrary time slice $t$ in the execution of a quantum program. There are $N$ logical qubits and $M$ hardware decoders ($N>M$). All decoding scheduling policies will have two components: The first will assign the decoders necessary for all critical decodes $C$ in the time slice $t$. The second will assign all the remaining $M-C$ hardware decoders to the $N-C$ qubits based on the scheduling policy used. We now discuss three decoder scheduling policies (all policies are illustrated in Fig. 6(b) – Fig. 6(d)):
@@ 4.2.1. Most Frequently Critically Decoded (MFD)
A logical qubit that consumes a significant number of $T$-states during the execution of a program would have a frequent requirement of critical decodes – leaving such a logical qubit undecoded for more than a few slices would make subsequent critical decodes take longer, thus slowing down computation. This motivates the MFD scheduling policy that prioritizes decoding of logical qubits that have numerous critical decodes in the future at any given time slice. The MFD policy will ensure that future critical decodes have a minimized number of undecoded syndromes for the qubits that have frequent critical decodes.
@@ 4.2.2. Round Robin (RR)
Derived from CPU scheduling policies used by operating systems, the RR policy does not prioritize any specific logical qubits – rather, it chooses $M-C$ qubits in a round-robin manner in every time slice to ensure fairness for all qubits in the system.
@@ 4.2.3. Minimize Longest Undecoded Sequence (MLS)
The longest undecoded sequence length at any given time slice is an indicator of how well the decoder scheduling policy is servicing all qubits in the system. We use this as a motivator for the MLS policy, which tries to minimize the longest undecoded sequence at every time slice. The MLS policy works as follows: at any time slice $t$, qubits are sorted on the basis of their current undecoded sequence lengths. Then, $M-C$ qubits with the largest undecoded sequence lengths are assigned hardware decoders.
@@ 4.3. Noise-Adaptive Decoder Scheduling
While control-flow instructions would necessitate runtime scheduling of decoders, events such as cosmic rays (McEwen2021, ) and leakage (Miao2023, ) can result in a temporary burst of errors for some physical qubits in the lattice that can impact some logical qubits. Scheduling after a control-flow instruction can be performed using any static scheduling policies for the program after the control-flow instruction. However, since the static scheduling policies do not account for the error-rate, spikes in errors due to cosmic rays and leakage cannot be factored at runtime without hardware support.
@@
Detecting a spike in the physical error rate will either require errors to be decoded or additional hardware modules to detect the spike. While error-correcting codes can tolerate temporary increases in the error-rate (Google2023, ), the increase in errors can result in longer decoding latencies since the decoding task becomes harder with more errors. If a logical qubit affected by these events is not scheduled for decoding immediately after the event, decoding it before applying a non-Clifford gate could take longer, thus delaying the operation and causing a slowdown. As shown in Fig. 7(a), an increase in the physical error rate results in a higher number of bit-flips (especially for larger code distances), which can be detected with simple components in the control hardware. Fig. 7(b) shows how additional flips can be detected and used to dynamically prioritize the decoding of an arbitrary patch $P$, which suffers from a temporary burst of errors. Note that the detection is different from decoding – we are merely predicting that there are more errors due to higher bit-flips in the syndromes.
@@ 4.4. Offloading to Software Decoders
Software decoders are slow and also have a higher variance in decoding latencies (Delfosse2023, ). However, when scheduling decoding tasks for logical qubits, software decoders can be leveraged to further reduce the undecoded sequence lengths. As shown in Fig. 7(c), some syndromes for a logical qubit can be offloaded to software while the hardware decoders are busy elsewhere. To prevent scheduled hardware decoding from being delayed, a buffer (three slices in this example) must be used to ensure that the software offloading completes before the next hardware decode.
5. Decoding for Distillation Factories
The decoder scheduling policies in the previous sections catered only to the decoding of algorithmic logical qubits (data logical qubits, magic state storage, ancillary logical qubits required for Lattice Surgery). In this section, we discuss decoding for distillation factories.
5.1. Distillation Factories
Magic state distillation factories generate few low-error logical qubits with non-Clifford states from many high-error logical qubits. Distillation factories run for very short periods at a time – this allows for smaller code distances to be used for creating the logical qubits for distillation (Litinski2019magic, ). As shown in Fig. 8, the error probability of a magic state is low enough to be useful even with $d=7$ ($d$ refers to $d_{X}$ in (Litinski2019magic, )).
Figure 8. Output error probabilities of magic states after 15-1 distillation for different $d$ (Litinski2019magic, ) ($p=10^{-4}$).
@@ 5.1. Distillation Factories
[fig]
5.2. Using Fast, Low-Footprint Decoders
Smaller code distances for distillation factories provide two main benefits: the number of physical qubits required are much lower, and more importantly, both hardware (Vittal2023, ; Smith2023, ; Delfosse2020, ; Caune2023, ) and software (Delfosse2023, ) decoders are faster and less complex. For example, LUT based decoders (Das2022lilliput, ) have been shown to be effective up to $d=5$ without requiring significant hardware resources. Predecoders (Delfosse2020, ; Ravi2023, ; Smith2023, ) reduce the complexity and decoding effort required for lower code distances as well. Compared to algorithmic logical qubits, which require a large code distance to survive millions of error correction cycles (Beverland2022, ; Blunt2024, ), the decoding requirements of distillation factories are far more relaxed, which reduces the hardware resource requirements as well.
The decoding overhead of magic state distillation is significantly lower than algorithmic logical qubits – it can thus leverage lightweight decoders, considerably reducing the hardware cost for distillation factories.
@@ 6. Methodology
We now describe the methodology used to evaluate different decoder scheduling policies and for estimating classical resources required for executing workloads on a Surface Code error-corrected quantum computer.
@@ 6.1. Compiler
We use the Lattice Surgery Compiler (LSC) (watkins2023high, ) to generate Intermediate Representations (IR) of workloads that can be executed on an error-corrected quantum computer using the Surface Code with lattice Surgery. LSC can generate IR that denote Lattice Surgery instructions from the QASM (qasm, ) representation of a workload. LSC handles mapping and routing based on the layout provided to the compiler. We configure LSC to use a ‘wave’ scheduling that maximizes the number of concurrent instructions executed in every time slice. LSC also uses Gridsynth (gridsynth, ) to deal with arbitrary rotations. However, since it is still under development, LSC has some limitations:
@@
  • LSC abstracts away distillation factories, only magic state storage sites are considered.
@@ 6.2. Simulation Framework
Using the IR generated by LSC, we build a framework that can parse the IR and determine the critical decodes in every slice, generate a timeline of all operations, and assign decoders to all logical qubits depending on the scheduling policy. In case the number of critical decodes in a particular slice are more than the number of hardware decoders configured, decoder-resources can rewrite the IR to defer critical decodes to the next slice (potentially increasing the execution time of the program).
Layouts: We use three layouts for our evaluations – Fast and Compact layouts (Litinski2019, ), and the EDPC layout (Beverland2022edpc, ). The Compact layout uses the fewest logical qubits and, due to a single routing lane, allows only one magic state to be consumed per time slice – we thus use it only to compare total execution times with the Fast and EDPC layouts.
Benchmarks: We use benchmarks from MQT Bench (quetschlich2023mqtbench, ) and QASMBench (Li2023, ). We use shor-15, a chemistry workload gndstate-14, a NISQ workload qaoa-14444Used for its arbitrary rotations and similarity to chemistry workloads., random, wstate, and arithmetic workloads adder-28, multiplier-45, Quantum Fourier Transform qft-20 – which can be used as building blocks for other algorithms.
Other Software: Stim (Gidney2021, ) was used for simulating stabilizer circuits to generate syndromes and error rates. Azure QRE (Beverland2022, ) was used for resource estimations.

### v2

@@ 3.3. Concurrency and Delayed Decoding
Critical decodes must be serviced during the execution of a program to avoid the increase in memory requirements and processing times discussed above. If there are many concurrent critical decodes that occur frequently during the execution of a program, an appropriate number of decoders will be needed to process these critical decodes. The level of concurrency depends entirely on the layout used to build the quantum computer – layouts such as the Compact (Litinski, 2019a, p. 7), Fast (Litinski, 2019a, p. 9), and the Edge-Disjoint Path (EDPC) compilation (Beverland et al., 2022a) are some proposed layouts that can be used to build fault-tolerant quantum computers using the Surface Code. Layouts determine the number of physical qubits required – for example, the Compact layout requires the fewest physical qubits since there is a single routing lane at the cost of completely serializing operations. The Fast and EDPC layouts allow more concurrency at the cost of more physical qubits.
What is the average level of concurrency when executing a quantum program on a fault-tolerant quantum computer? Figure 4(c) shows a histogram of the number of critical decodes [^2] For the Surface Code, we assume every logical qubit requires two decoders – one each for the $X$ and $Z$ observables. for select workloads generated by the Lattice Surgery Compiler (Watkins et al., 2024). This histogram shows that the peak concurrency is attained very infrequently. This implies that most logical qubits function as memory qubits or execute Clifford gates more often than $T$-gates, and this can allow some qubits to not be decoded in real-time.
Quantum programs are serial in terms of $T$-gates applied – not every qubit requires constant access to a fast decoder.
@@ 3.4. Goal: Make Classical Processing Efficient
Provisioning a hardware decoder for every logical qubit in the system can be resource intensive. Figure 4(d) shows an estimate of the total number of decoders required for all logical qubits (=algorithmic, ancillary, magic state storage, and magic state distillation logical qubits) when using the EDPC layout for different workloads. These estimates were obtained by using Lattice Surgery Compiler (Leblond et al., 2024; Watkins et al., 2024). Note that the total hardware requirement will be significantly higher because of control and readout components. Having shown how the syndrome processing rate is crucial in ensuring that computation does not require excessive memory and time and how quantum programs are inherently serial in terms of critical decodes, the question we seek to answer in this work is–
How can we minimize the use of hardware decoders and lower classical processing costs without sacrificing performance and reliability?
4. VQD: Virtual Quantum Decoding
We now discuss decoder virtualization, where there are fewer hardware decoders than logical qubits in the system, and how decoding can be scheduled optimally.
4.1. Working with Fewer Hardware Decoders
Reducing the number of available decoders implies that qubits will share hardware resources, resulting in time-division multiplexing of decoder instances among logical qubits. Figure 5 shows how compared to a system with decoders for every logical qubit in the system ($N$ qubits, $N$ decoders), a system with fewer ($M$, $N>M$) decoders will require resources to be shared with time.
Before we discuss different decoder scheduling policies, it is important to understand the time granularity at which any scheduling policy will operate on. Since logical operations (Clifford or non-Clifford) in a Surface Code error-corrected quantum computer will require at least $d$ rounds before the next operations, we define a slice (Watkins et al., 2024) as the smallest time step between logical operations that a decoder scheduling policy can work on. Every slice consists of $d$ rounds of syndrome measurements, thus making the scheduling policy agnostic of the actual code distance used.
@@ 4.1. Working with Fewer Hardware Decoders
4.2. Decoder Scheduling Policies
Static decoder scheduling refers to decoder scheduling that can be performed at compile time. Since most quantum programs do not have any control-flow instructions, scheduling can be performed statically. Scheduling decoders is similar to CPU scheduling performed by all operating systems today, where the number of processes is more than the number of available processor cores (Liu and Layland, 1973).
Longest Undecoded Sequence: To quantify the fairness of a decoder scheduling policy, we use ‘Longest Undecoded Sequence’, which measures how well the decoders are servicing all logical qubits. A large undecoded sequence length implies that a qubit has been left undecoded for a long time – increasing the memory consumed to store undecoded syndromes. Figure 6(a) shows an example of determining the longest undecoded sequence length.
Consider an arbitrary time slice $t$ in the execution of a quantum program. There are $N$ logical qubits and $M$ hardware decoders ($N>M$). All decoding scheduling policies will have two components: The first will assign the decoders necessary for all critical decodes $C$ in the time slice $t$. The second will assign all the remaining $M-C$ hardware decoders to the $N-C$ qubits based on the scheduling policy used. We now discuss three decoder scheduling policies (all policies are illustrated in Figure 6(b) – Figure 6(d)):
@@ 4.2.1. Most Frequently Critically Decoded (MFD)
A logical qubit that consumes a significant number of $T$-states during the execution of a program would have a frequent requirement of critical decodes – leaving such a logical qubit undecoded for more than a few slices would make subsequent critical decodes take longer, thus slowing down computation. This motivates the MFD scheduling policy that prioritizes decoding of logical qubits that have numerous critical decodes in the future at any given time slice (Figure 6(b)). The MFD policy will ensure that future critical decodes have a minimized number of undecoded syndromes for the qubits that have frequent critical decodes.
@@ 4.2.2. Round Robin (RR)
Derived from CPU scheduling policies used by operating systems, the RR policy does not prioritize any specific logical qubits – rather, it chooses $M-C$ qubits in a round-robin manner (Figure 6(c)) in every time slice to ensure fairness for all qubits in the system.
@@ 4.2.3. Minimize Longest Undecoded Sequence (MLS)
The longest undecoded sequence length at any given time slice is an indicator of how well the decoder scheduling policy is servicing all qubits in the system. We use this as a motivator for the MLS policy, which tries to minimize the longest undecoded sequence at every time slice (Figure 6(d)). The MLS policy works as follows: at any time slice $t$, qubits are sorted on the basis of their current undecoded sequence lengths. Then, $M-C$ qubits with the largest undecoded sequence lengths are assigned hardware decoders.
@@ 4.3. Noise-Adaptive Decoder Scheduling
While control-flow instructions would necessitate runtime scheduling of decoders, events such as cosmic rays (McEwen et al., 2021) and leakage (Miao et al., 2023) can result in a temporary burst of errors for some physical qubits in the lattice that can impact some logical qubits. Detecting a spike in the physical error rate will either require errors to be decoded or additional hardware modules to detect the spike. As shown in Figure 7(a), an increase in the physical error rate results in a higher number of bit-flips (especially for larger code distances), which can be detected with simple components in the control hardware. Figure 7(b) shows how additional flips can be detected and used to dynamically prioritize the decoding of an arbitrary patch $P$, which suffers from a temporary burst of errors. Note that the detection is different from decoding – we are merely predicting that there are more errors due to higher bit-flips in the syndromes. Unlike prior work (Suzuki et al., 2022), this mechanism only prioritizes the decoding of logical qubits affected by error bursts in a virtualized decoder environment. However, this prioritization can result in reduced efficacy of virtualization, due to fewer decoders being available for scheduling during error burst events. We evaluate the overheads of such error bursts for virtualization in Section 7.
@@ 4.4. Offloading to Software Decoders
Software decoders are slow and also have a higher variance in decoding latencies (Delfosse et al., 2023). However, when scheduling decoding tasks for logical qubits, software decoders can be leveraged to further reduce the undecoded sequence lengths. As shown in Figure 7(c), some syndromes for a logical qubit can be offloaded to software while the hardware decoders are busy elsewhere. To prevent scheduled hardware decoding from being delayed, a buffer (three slices in this example) must be used to ensure that the software offloading completes before the next hardware decode.
4.5. Decoding for Distillation Factories
The decoder scheduling policies in the previous sections catered only to the decoding of algorithmic logical qubits (data logical qubits, magic state storage, ancillary logical qubits required for Lattice Surgery). Magic state distillation factories generate few low-error logical qubits with non-Clifford states from many high-error logical qubits. Distillation factories run for very short periods at a time – this allows for smaller code distances to be used for creating the logical qubits for distillation (Litinski, 2019b). Smaller code distances for distillation factories provide two main benefits: the number of physical qubits required are much lower, and more importantly, both hardware (Vittal et al., 2023; Smith et al., 2023; Delfosse, 2020; Caune et al., 2023) and software (Delfosse et al., 2023) decoders are faster and less complex. For example, LUT based decoders (Das et al., 2022a) have been shown to be effective up to $d=5$ without requiring significant hardware resources. Predecoders (Delfosse, 2020; Ravi et al., 2023; Smith et al., 2023) reduce the complexity and decoding effort required for lower code distances as well. Compared to algorithmic logical qubits, which require a large code distance to survive millions of error correction cycles (Beverland et al., 2022b; Blunt et al., 2024), the decoding requirements of distillation factories are far more relaxed, which reduces the hardware resource requirements as well.
Magic state distillation has lower hardware requirements for decoding due to lower code distances involved – and virtualization can further reduce them.
4.6. Impact of Decoding Latencies
So far, we have considered decoder virtualization without considering the decoder latency. However, the former is heavily impacted by the latter. Consider a scenario where a logical qubit has $R$ undecoded rounds prior to a critical/scheduled decode. Now, the decoder will take a finite time to process these $R$ rounds worth of syndromes based on its per round decoding latency $T_{proc}$. However, during the time the initial $R$ rounds of syndromes are processed, syndromes are continuously generated based on the syndrome generation time $T_{gen}$. Thus, at any time $t$, the total rounds processed are $P(t)=t/T_{proc}$, the syndrome rounds generated are $G(t)=t/T_{gen}$, and so, the outstanding rounds to be processed can be defined as $S(t)=R+G(t)-P(t)$. The total time taken for the decoder to catch up is thus $t_{total}$, with $S(t_{total})=0$, and hence:
(1) $t_{total}=\frac{R}{T_{proc}^{-1}-T_{gen}^{-1}};(iff~{}~{}T_{proc}<T_{gen})$
Figure 8. (a) Slowdown in the processing of outstanding syndromes for different normalized decoder latencies $t_{D}=T_{proc}/T_{gen}$. Slower decoders require more time (and thus more rounds) to catch up with the present syndrome; (b) Differences in the total decoding task for different initial undecoded rounds – a $t_{D}=0.99$ increases the number of rounds to be decoded by a 100$\times$.
@@ 4.6. Impact of Decoding Latencies
[some figure]
$T_{proc}$ has a considerable effect on the efficacy of decoder virtualization, since a slower processing rate would mean that a huge number of additional rounds will be required for the decoder to be able to catch up to the present syndrome being generated. This can be seen from Figure 8(a), where the decoder latency is normalized by the syndrome generation time $t_{D}=T_{proc}/T_{gen}$ and the slowdown is defined as the ratio of the total time taken to process $R$ undecoded rounds (Equation 1) and the ideal time when extra syndromes are added to the decoding task ($t_{total}^{\prime}=R/T_{proc},slowdown=t_{total}/t_{total}^{\prime}$). This slowdown is agnostic of the actual implementation of the decoder – techniques like parallel window decoding (Skoric et al., 2023) will only help reduce the value of $t_{D}$. Figure 8(a) shows the total decoding task ($P(t_{total})$), which represents the total syndromes that need to be processed for an initial $R$ undecoded rounds. From Figure 8(a), we see that the closer the decoding processing rate is to the syndrome generation rate ($t_{D}=0.99$), the harder virtualization will be, since even $R=10$ will require $10^{4}$ rounds before the decoder can catch up. This is a crucial limitation for decoder virtualization, and it also suggests that faster, more accurate decoders will be more desirable than fast decoders which have a smaller hardware footprint at the cost of accuracy. Figure 8(b) shows an example for cases where there are 10/100 undecoded rounds before the decoder is invoked for a logical qubit – a slower decoder will require thousands of extra rounds to catch up compared to a faster decoder! Running so many additional rounds will slow down the system, reduce the error budget per round of computation, exacerbating the logical error rates. In Section 7, we will show how slow decoders associated with qLPDC codes are not amenable for virtualization, due to the slowdown they incur for even modest undecoded sequence lengths. This represents a limit of virtualization, and motivates further research in developing fast, accurate decoders for all QEC codes.
Figure 9. System architecture for virtualized decoding.
[…]
[some figure]
5. System Architecture with Virtual Decoders
Enabling the virtualization of decoders in an FTQC system requires (i) memory to store syndromes of logical qubits that are not being decoded at any time, and (ii) a mechanism for selecting the qubits to be decoded by a specific decoder at any time. Figure 9 shows the system architecture that can be used for enabled virtualized decoders. The Arbiter is a key component that routes syndromes from the readout system to the decoders. Since there will be fewer decoders than logical qubits, syndromes will be routed based on the static decoding schedule prepared during compile time. To allow decoders to access the syndromes of a specific logical qubit at any given point in time, the arbiter attaches a tag to the syndromes before routing them to the appropriate decoder. This tag quickly identifies the logical qubit to which a syndrome belongs to. Every decoder consists of a syndrome memory, and the decoding schedule can be used by the control software to access the syndromes of any logical qubit from the syndrome memory.
Software offloading can be easily integrated in the design by enabling the syndrome memory to be read from the control software. Note that this will require high-speed links between the control software and the decoder modules, such as the low-latency ethernet used in recent experimental demonstrations (Acharya et al., 2024). The control software can specify the tag and the syndromes to be read from the syndrome memory. This will however require fusing the software decoding results with the hardware decoding results. Fusion of separate decoding tasks has also been demonstrated recently by Google (Acharya et al., 2024). To enable dynamic decoder scheduling, the control software need only trigger decoding for the necessary logical qubit(s) by amending the decoding schedule. Since this could trigger the decoding of any arbitrary logical qubit, the syndrome memories must support random accesses. No other changes to the Arbiter are necessary, since syndromes can be routed to the same memory regardless of dynamic scheduling that could occur later in time.
Figure 10. Compilation framework – the rewrite pass changes the IR schedule based on the number of available decoders.
[…]
[some figure]
@@ 6. Methodology
We now describe the methodology used to evaluate different decoder scheduling policies and classical processing requirements.
@@ 6.1. Compiler
We use the Lattice Surgery Compiler (LSC) (Watkins et al., 2024) to generate Lattice Surgery Instructions (LLI) Intermediate Representations (IR) of workloads that can be executed on an error-corrected quantum computer using the Surface Code with lattice Surgery. LSC can generate IR that denote Lattice Surgery instructions from the QASM (Cross et al., 2017) representation of a workload. The Lattice Surgery instructions generated by LSC are a combination of Clifford operations and multi-body measurements used for implementing Pauli-Product Rotations (PPR) (Litinski, 2019a). LSC handles mapping and routing based on the layout provided to the compiler, as shown in Figure 10. We configure LSC to use a ‘wave’ scheduling that maximizes the number of concurrent instructions executed in every time slice. LSC uses Gridsynth (Ross and Selinger, 2014) to deal with arbitrary rotations. However, since it is still under development, LSC has some limitations:
@@
  • LSC treats distillation factories as black boxes.
@@ 6.2. Simulation Framework
Using the IR generated by LSC, we build a framework [^3] https://anonymous.4open.science/r/decoder-resources-5EC4/ that can parse the IR and determine the critical decodes in every slice, generate a timeline of all operations, and assign decoders to all logical qubits depending on the scheduling policy. In case the number of critical decodes in a particular slice are more than the number of hardware decoders configured, decoder-resources can rewrite the IR to defer critical decodes to the next slice.
Layouts: We use three layouts for our evaluations – Fast, Compact (Litinski, 2019a), and the EDPC (Beverland et al., 2022a) layouts. The Compact layout uses the fewest logical qubits and, due to a single routing lane, allows only one magic state to be consumed per time slice – we thus use it only to compare total execution times with the Fast and EDPC layouts.
Benchmarks: We use benchmarks from (Quetschlich et al., 2023) and (Li et al., 2023). We use shor-15, phase estimation qpe-50, Quantum Fourier Transform qft-20, arithmetic workloads adder-28, multiplier-45. We also use a neural network workload dnn-51, W state workloads wstate-40/60, random unitary random-40, and an Ising model ising-66.
Other Software: Stim (Gidney, 2021) was used for simulating stabilizer circuits to generate syndromes and error rates. Azure QRE (Beverland et al., 2022b) was used for resource estimations.